OpenStudy (anonymous):

let y=f(x)=(3x+7)/(x-2). find the inverse function.l

OpenStudy (amistre64):

solve for x as a function of y: x=f(y)

OpenStudy (amistre64):

y =(3x+7)/(x-2) ; multiply off the (x-2) y(x-2) = 3x+7 ; exand to get to the x part yx -2y = 3x + 7 ; -yx and -7 -2y-7 = 3x -yx ; factor out the x -2y-7 = x(3-y) ; divide out the (3-y) -2y-7 ----- = x ; simplify as wented 3-y